# Recursive Deep Inverse Reinforcement Learning

### v4

@@ A.3 RDIRL
A.4 Derivation of the recursive second-order Newton solution
In a similar fashion to Kalman filtering optimization process described in [Humpherys et al., 2012], we seek to determine optimal solution $\Theta_{N}^{*}=\{\theta^{*}(t_{0}),\dots,\theta^{*}(t_{N})\}$ using the second-order Newton method sequentially, which recursively finds $\Theta_{N}^{*}$ given $\Theta_{N-1}^{*}$. To do so, we start by breaking the optimization function (11) as follows:
@@ A.4 Derivation of the recursive second-order Newton solution
Next, we further divide (14) into the following form
@@
Our optimization approach consists of minimizing (16) then minimizing (15) given (16) and the minimizer $\hat{\Theta}_{i|i-1}$ of (16). We proceed by minimizing (16) with respect to $\Theta_{i}$ by finding $\Theta_{i}$ that drives the gradient of (16) to zero. By taking the gradient of Equation (16) with respect to $\Theta_{i}$ we obtain:
@@
Given (18) and (16), we proceed to minimize (15) using the second-order Newton update. We start by deriving the gradient of (15) as follows:
@@
Therefore at $\Theta_{i}=\hat{\Theta}_{i|i-1}$ Equation (19) becomes:
@@
The resulting optimal variable $\hat{\theta}(t_{i})\in\hat{\Theta}_{i}$ is given by Equation (12). The procedure is repeated until $t_{i}=t_{N}$.
A.5 Additional Experiments Results
A.5.1 Online Adaptation of competing methods
In this section, we compare our proposed approach, RDIRL, with online-adapted versions of GAIL, AIRL, and GCL. The online adaptation involves training each competing method using one expert demonstration at a time. Specifically, the loss function of each method is computed using a single observed expert sample at each time step, followed by an immediate update of the reward function neural network parameters. This process is repeated across the full episode of Nsteps.
As illustrated in Figure 4, our proposed method consistently outperforms the online-adapted baselines. Furthermore, the online adaptation does not significantly improve the performance of the original methods. In the case of Cartpole, it even leads to notable performance degradation and increased instability compared to both the original baselines (GAIL, AIRL, GCL) and our approach, as shown in Table 4. These results highlight the advantage of our recursive optimization framework in producing more stable and accurate reward functions over naive online adaptation.
@@ A.5.1 Online Adaptation of competing methods
GAIL $140.22\pm 8.75$ $-26.09\pm 58.04$
GCL $138.08\pm 13.83$ $-11.34\pm 64.05$
AIRL $142.96\pm 10.39$ $-27.13\pm 55.62$
GAIL-Online $99.38\pm 58.13$ $-5.39\pm 56.47$
GCL-Online $119.28\pm 50.95$ $4.52\pm 65.59$
AIRL-Online $113.48\pm 52.80$ $-6.76\pm 55.03$
RDIRL (ours) $\bm{149.02\pm 2.09}$ $\bm{47.16\pm 66.72}$

### v5

@@ B.3 RDIRL
B.4 Derivation of the recursive second-order Newton solution
In a similar fashion to Kalman filtering optimization process described in (Humpherys et al., 2012), we seek to determine optimal solution $\Theta_{N}^{*}=\{\theta^{*}(t_{0}),\dots,\theta^{*}(t_{N})\}$ using the second-order Newton method sequentially, which recursively finds $\Theta_{N}^{*}$ given $\Theta_{N-1}^{*}$. To do so, we start by breaking the optimization function (11) as follows:
@@ B.4 Derivation of the recursive second-order Newton solution
Next, we further divide equation 14 into the following form
@@
Our optimization approach consists of minimizing equation 16 then minimizing equation 15 given equation 16 and the minimizer $\hat{\Theta}_{i|i-1}$ of equation 16. We proceed by minimizing equation 16 with respect to $\Theta_{i}$ by finding $\Theta_{i}$ that drives the gradient of equation 16 to zero. By taking the gradient of equation 16 with respect to $\Theta_{i}$ we obtain:
@@
Given equation 18 and equation 16, we proceed to minimize equation 15 using the second-order Newton update. We start by deriving the gradient of equation 15 as follows:
@@
Therefore, at $\Theta_{i}=\hat{\Theta}_{i|i-1}$, equation 19 becomes:
@@
The resulting optimal variable $\hat{\theta}(t_{i})\in\hat{\Theta}_{i}$ is given by equation 12. The procedure is repeated until $t_{i}=t_{N}$.
B.5 Additional Experiments Results
B.5.1 Online Adaptation of competing methods
In this section, we compare our proposed approach, RDIRL, with online-adapted versions of GAIL, AIRL,ML-IRL, and GCL. The online adaptation involves training each competing method using one expert demonstration at a time. Specifically, the loss function of each method is computed using a single observed expert sample at each time step, followed by an immediate update of the reward function neural network parameters. This process is repeated across the full episode of Nsteps.
As illustrated in Figure 4, our proposed method consistently outperforms the online-adapted baselines. Furthermore, the online adaptation does not significantly improve the performance of the original methods. In the case of Cartpole, it even leads to notable performance degradation and increased instability compared to both the original baselines (GAIL, AIRL,ML-IRL, GCL) and our approach, as shown in Table 4. These results highlight the advantage of our recursive optimization framework in producing more stable and accurate reward functions over naive online adaptation.
@@ B.5.1 Online Adaptation of competing methods
GAIL $0.934\pm 0.058$ $0.521\pm 1.15$
GCL $0.92\pm 0.09$ $0.226\pm 1.27$
AIRL $0.953\pm 0.069$ $0.54\pm 1.11$
GAIL-Online $74\pm 0.29 0.0$$02\pm 0.51$
GCL-Online $0.84\pm 0.25$ $0.1\pm 0.53$
AIRL-Online $81\pm 0.26 0.0$$01\pm 0.49$
ML-IRL-Online $0.49\pm 0.29$ $0.14\pm 0.75$
RDIRL (ours) $\bm{0.99\pm 0.13}$ $\bm{0.49\pm 0.59}$